OpenStudy (aivantettet26):

find the equations of the tangent lines to 9x^2 + 16y^2 = 52 that are parallel to the line 9x-8y =1

OpenStudy (amistre64):

define your dy/dx

OpenStudy (aivantettet26):

81/128

OpenStudy (aivantettet26):

slope = 9/16

OpenStudy (aivantettet26):

correction ! slope = 81/128

OpenStudy (amistre64):

9x^2 + 16y^2 = 52 9x + 16y y' = 0 y' = -9x/16y

OpenStudy (amistre64):

for this to be parallel to the given line; we want it to equal: 9x-8y =1 9-8y' =0 y' = 9/8

OpenStudy (aivantettet26):

since slope = y/x meaning that the points are (8 , 9)?

OpenStudy (amistre64):

\[\frac{9x}{16y}=-\frac{9}{8}\]solve for x,y

OpenStudy (amistre64):

we have to equate the derivatives to determine the (x,y) points

OpenStudy (aivantettet26):

ohhhh... okay okay

OpenStudy (amistre64):

i misplaced a negative, but i think it amounts to the same thing :)

OpenStudy (aivantettet26):

72x = 144y

OpenStudy (aivantettet26):

x = 2?

OpenStudy (amistre64):

\[-\frac{9x}{16y}=\frac{9}{8}\] \[-\frac{9x}{1}=\frac{9~2y}{1}\] \[-x=2y~:~y=-\frac12x\]

OpenStudy (anonymous):

no x = 2y

OpenStudy (aivantettet26):

oops! i forgot y XD

OpenStudy (anonymous):

and actually its x = -2y , I think ..

OpenStudy (anonymous):

Or we can just look at the great work that @amistre64 has done.

OpenStudy (amistre64):

9x^2 + 16y^2 = 52; x=-2y is fine too 94y^2 + 16y^2 = 52

OpenStudy (aivantettet26):

yes. negative sign too since it was misplaced hahaha

OpenStudy (amistre64):

solving for the x,y is just a matter of plugging in .... then you can apply the slope to those points

OpenStudy (amistre64):

that not 94y^2, its 9*4y^2 :)

OpenStudy (aivantettet26):

y = 1?

OpenStudy (amistre64):

y = 1 or -1 yes

OpenStudy (aivantettet26):

since y = 1 x = 2y x = -2(1) x = -2

OpenStudy (amistre64):

good, x = 2 or -2 i believe

OpenStudy (amistre64):

you have 4 points, and can prolly narrow it down to the correct 2 since the original i believe is an ellipse

OpenStudy (aivantettet26):

9x + 8y =-26

OpenStudy (anonymous):

Actually you only have 2 points. It can only be y = -1 or 1. When y = -1 x = 2 and when y = 1 x = -2 so the two points are (-2, 1) and (2,-1)

OpenStudy (amistre64):

(-2,1) and (2,-1) are lovely

OpenStudy (aivantettet26):

wow. that is a really neat graph!

OpenStudy (amistre64):

it took me weeks to make it lol

OpenStudy (anonymous):

Sweet wolfram curve @amistre64 . I need to become more adept at using that site.

OpenStudy (anonymous):

lol yeeeeaaaaahhh

OpenStudy (aivantettet26):

thank you for all of your efforts. both of you!

OpenStudy (anonymous):

Credit goes to @amistre64 he/she carried the majority of the work load!

OpenStudy (anonymous):

I just made comments from the cheap seats.

OpenStudy (amistre64):

its hard to seperate the extra chatter from the voices in my head :)